# Target Aggregate Data Adjustment Method for Transportability Analysis Utilizing Summary-Level Data from the Target Population

### v1

@@ 3.2 Censoring Weights: Inverse Probability of Censoring Weight
In longitudinal studies, participants are lost to follow-up for various reasons that may be related to the treatment or competing events. This loss to follow-up can result in systematic differences between those who remain in the study and those who do not, across treatment groups. While RCTs generally eliminate confounding through randomization, censoring can still introduce bias if the likelihood of being censored is related to the outcome the participant would have experienced in the absence of censoring. To address this issue, censoring weights are applied during the estimation process, aiming to recreate a hypothetical population that would have been observed without censoring. TADA aims to apply inverse probability censoring weights (IPCW) to account for bias resulting from censoring for the time-varying outcome in the source population.
Let $A=1$ represent the presence of treatment and $A=0$ represent the absence of treatment. Let $c$ be the censoring indicator where $c=1$ represents being censored and $c=0$ represents not being censored. Let X represent the baseline covariates related to censoring and the outcome. The standard inverse probability of censoring weight $w^{\textbf{c}}$ is the inverse of the conditional probability of not being censored given the baseline covariates and treatment status:
$w^{c}=\frac{1}{\Pr(c=0\mid\textbf{X},A)}.$
In an RCT, where patients are allocated to two parallel arms, the primary endpoint, such as overall survival, might be right-censored. In this case, at any given time $t$, the censoring weights are defined as the inverse of the probability that an individual remained on the originally randomized treatment (i.e., uncensored) until time $t$. The weights for each individual up to the censoring time can be defined as:
$w_{i}^{\textbf{c}}=\frac{1}{P(C_{i}>T_{i}\mid\textbf{X}=\textbf{x}_{i})},$
where $C_{i}$ is the time to censoring for individual $i$ and $T_{i}$ is the time to event for individual $i$.
To address the issue of unstable weights due to possibly extreme probability as denominators, stabilized weights are often used with the advantages of less variable and less skewed, as suggested by Robins 31. The stabilized censoring weights are defined as:
$sw_{i}^{\textbf{c}}=\frac{\bar{P}(C_{i}>T_{i})}{P(C_{i}>T_{i}\mid\textbf{X}=\textbf{x}_{i})},$
where the weight denominators remain constant, the numerators reflect the overall survival probability (the average probability of remaining uncensored).
We utilize the Weibull distribution to model both event times and censoring times, implying that the baseline hazard functions are time-varying 32, 33. The hazard function of the Weibull distribution is defined as:
@@
where $\lambda$ is the scale parameter, $\alpha$ is the shape parameter, and $t$ represents time. When $\alpha\neq 1$, the hazard function varies with time: If $\alpha>1$, the hazard increases over time, If $\alpha<1$, the hazard decreases over time.
The cumulative hazard function is:
$H(t)=\lambda t^{\alpha},$
and the survival function is:
$S(t)=\exp\left\{-H(t)\right\}=\exp\left(-\lambda t^{\alpha}\right).$
We use the Cox PH model 34, 35 to model the relationship between censoring times and covariates:
$h_{c}(t\mid\textbf{X}_{i})=h_{c0}(t)\exp\left(\beta_{c}^{\top}\textbf{X}_{i}\right),$
where $h_{c}(t\mid X_{i})$ is the hazard function for censoring for individual $i$ at time $t$, $h_{c0}(t)$ is the time-varying baseline hazard function, $\beta_{c}$ is the vector of regression coefficients, and $\textbf{X}_{i}$ is the vector of covariates.
We estimate the baseline cumulative hazard function $H_{c0}(t)$ using the Breslow estimator 36:
[…]
For each individual $i$, we calculate the linear predictor:
$\eta_{i}=\beta_{c}^{\top}\textbf{X}_{i}.$
The probability that individual $i$ is uncensored at their observed time $t_{i}$ is:
$S_{c}(t_{i}\mid\textbf{X}_{i})=\exp\left\{-H_{c0}(t_{i})\exp\left(\eta_{i}\right)\right\}.$
The censoring weight $w_{i}$ for individual $i$ is defined as the inverse of their uncensored probability:
$w^{\textbf{c}}_{i}=\frac{1}{S_{c}(t_{i}\mid\textbf{X}_{i})},$
and the final stabilized censoring weights $sw_{i}^{\textbf{c}}$ are then used in TADA to adjust for bias due to censoring.
3.3 Final Weights: A Two-Staged Scheme
We introduce the final weights $w_{i}^{\textbf{final}}$ for each individual as the product of participation weights $w_{i}^{\textbf{MoM}}$ and the stabilized censoring weights $sw_{i}^{\textbf{c}}$ for all covariates of interest selected for population matching adjustment as follows:
$w_{i}^{\textbf{final}}=w_{i}^{\textbf{MoM}}\times sw_{i}^{\textbf{c}}.$
To mitigate the negative impact of extreme weights on estimation, we apply truncation at the 99% quantile. The final weights adjust for both the bias from censoring and the differences in effect modifier distributions between the source and target populations. This combined weighting approach ensures unbiased treatment effect estimates concerning both censoring and population differences. The weighted source population should closely resemble the target population, demonstrating population balance and supporting the treatment effect estimates.
In TADA, we employ a marginal structural model (MSM) to estimate the causal effect of treatment on survival outcomes. Since our weights are solely based on baseline covariates and do not involve any time-dependent components, we use a simplified MSM suitable for this context 31, 37. The MSM is specified as a weighted Cox PH model 34:
$\lambda_{a}(t\mid A_{i})=h_{0}(t)\exp(\beta A_{i}),$
where $h(t\mid A_{i})$ is the hazard function at time $t$ for individual $i$ with treatment assignment $A_{i}$, $h_{0}(t)$ is the baseline hazard function, and $\beta$ represents the log hazard ratio of the treatment effect.
We estimate $\beta$ by fitting the weighted Cox model using the final weights $w_{i}^{\textbf{final}}$ obtained from our two-stage weighting scheme as detailed in Sections 3.1 to 3.3. By applying these final weights in the Cox model we create a pseudo-population in which: the treatment is independent of baseline covariates; the bias due to censoring is reduced; and balance in effect modifiers between the source and target populations is achieved. This allows us to estimate the transported treatment effect on the target population using aggregate-level data.
The weighted partial likelihood function for the Cox model 38 is given by:
$L(\beta)=\prod_{i=1}^{n_{s}}\left[\frac{\exp(\beta A_{i})}{\sum_{j\in R(t_{i})}w_{j}^{\textbf{final}}\exp(\beta A_{j})}\right]^{\delta_{i}w_{i}^{\textbf{final}}},$
where $n_{s}$ is the number of individuals in the source sample, $\delta_{i}$ is the event indicator for individual $i$, and $R(t_{i})$ is the risk set at time $t_{i}$. Maximizing this weighted likelihood function provides an estimate of $\beta$, the log hazard ratio, which quantifies the causal effect of the treatment in the target population. We are estimating the variance of $\beta$ using the bootstrap method.
By integrating the MSM into TADA, the treatment effect estimator is unbiased concerning both censoring and population imbalance. The weighted source sample is adjusted to closely resemble the target population in terms of aggregate-level summaries, thereby enhancing the transportability of causal inferences.
4 Simulation Study
The simulation study applies the ADEMP framework proposed by Morris et al. 39 to comprehensively evaluate the effectiveness of TADA in transportability analysis using AgD for a target population, specifically examining its performance in adjusting for censoring. Our primary goal (Aim) is to assess the estimated causal effects of treatment on overall survival across different RCT scenarios, with and without censoring adjustments made via TADA. Detailed descriptions of the Data-generating mechanisms, Estimands, Methods, and Performance measures are provided in the following subsections.
4.1 Data-Generating Mechanism
In each simulation replicate, we generate three datasets named studyData, targetData, and targetData_Popula-tion_IPD respectively: studyData contains simulated individual-level data from the source population, including treatment status, covariates, event times, censoring times, observed times, and indicators for observed and censored events. targetData provides aggregate summary statistics (proportions, means, and standard deviations of each covariate) for the target population, matching those in studyData. targetData_Population_IPD consists of pseudo individual-level data for the target population, including covariates, treatment status, and event times.
@@ 4.1 Data-Generating Mechanism
To generate studyData, we first simulate a large source population of 50,000 individuals and then randomly sampled 200 individuals for the final dataset. Each individual has three baseline covariates: $X_{1}$ and $X_{2}$ are binary variables generated from binomial distributions with parameters 0.45 and 0.65, respectively. $X_{3}$ is a continuous variable generated from a standard normal distribution (mean 0, variance 1). Treatment status is assigned randomly to simulate an RCT scenario, with an allocation rate of 1:1.
@@
Event times are generated from a Weibull distribution, incorporating the effects of covariates and treatment through the linear predictor $\eta_{\text{event},i}$.
@@
Event times $T_{i}$ are simulated to model time-to-event data with a time-varying baseline hazard function using the inverse cumulative distribution function (CDF) of Weibull distribution:
@@
In studyData, we generate censoring time for each individual following a Weibull distribution with the impact of both treatment and covariates. A linear predictor $\eta_{\text{censor},i}$ is constructed for each individual to incorporate the effects of covariates on the hazard function.
@@
where $\beta_{X_{1}}^{c}=-0.2$, $\beta_{X_{2}}^{c}=0.4$ and $\beta_{X_{3}}^{c}=-0.1$ as the main effect of baseline covariates, $\beta_{\text{trt}}^{c}=0.25$ as the treatment main effect and $\beta_{0}$ as the intercept. We choose various candidates of $\beta_{0}$ to directly adjust the overall censoring proportion among studyData. We set $\beta_{0}$ as 2.5, 3.3, 3.7 and 4.3 respectively to obtain studyData with around 20%, 30%, 40% and 50% overall censoring proportion while keeping other coefficients being consistent.
@@
For each individual, the observed time ($U_{i}$) is determined by:
@@
The event indicator $\Delta_{i}$ is defined as:
@@
Similarly, we generate targetData by simulating a large target population of 200,000 individuals and then randomly sampling 1,000 individuals for the final dataset. Each individual has three baseline covariates as same as in the source population ($X_{1}$, $X_{2}$, and $X_{3}$) with different distribution characteristics: $X_{1}$ is binary and generated from a binomial distribution with parameter 0.35. $X_{2}$ is binary and generated from a binomial distribution with parameter 0.55. $X_{3}$ is continuous and generated from a normal distribution with a mean of 0.21 and a variance of 1.5. Since targetData is an aggregate-level dataset, we summarize the proportions of $X_{1}$ and $X_{2}$, and the mean and standard deviation of $X_{3}$ as the final data.
The individual-level covariate data for targetData_Population_IPD are generated concurrently before aggregation. To obtain the “true” treatment effect in the target population for benchmarking, we assign treatment status and generate event times for each individual in targetData_Population_IPD. Treatment status is randomly assigned with a 50% probability to simulate an RCT scenario. We use the same Weibull distribution parameters ($\alpha_{\text{event}}$, $\lambda_{\text{event}}$), the same linear predictor structure $\eta_{\text{event},i}$ and coefficients as in studyData, applying them to targetData_Population_IPD to ensure consistency.
4.2 Estimands, Targets and Methods
@@
  • Adjusting censoring: In this case, TADA adjusts the bias resulting from the presence of censoring via obtaining the IPCW introduced in Section 3.1, as well as balancing the source and target population with the participation weights introduced in Section 3.2. The final weights are as introduced in Section 3.3 comprised of both weights.
  • Ignoring censoring: In this case, TADA ignores the impact of the presence of censoring and only balances the source and target population with the participation weights introduced in Section 3.2. The final weights are equivalent to the participation weights.
For both strategies, we estimate average marginal hazard ratios using a Cox PH model based on multiple replicates. Meanwhile, we apply the Cox PH model on targetData_Population_IPD to estimate average marginal hazard ratios as the pseudo “true HR” of the target population as the comparison benchmark.
@@ 4.3 Performance Measures
The performance measures of interest are the bias of the estimators calculated using the above methods and their coverage of the 95% confidence intervals compared to the analysis in the IPD target data. Let $\hat{\gamma}_{\text{cen,i}}$ be the estimated marginal hazard ratios for the $i=1,2,\cdots,N$-th simulated replicate when considering the adjustment of censoring. Let $\hat{\gamma}_{\text{non-cen,i}}$ be the estimated marginal hazard ratios for the $i=1,2,\cdots,N$-th simulated replicate when ignoring the adjustment of censoring. With the pseudo “true HR” of target population being $\bar{\gamma}$, the bias of two cases aforementioned are estimated with
$\widehat{\text{Bias}}_{\text{cen}}$ $=\frac{1}{N}\sum_{i=1}^{N}\hat{\gamma}_{\text{cen,i}}-\bar{\gamma},$
$\widehat{\text{Bias}}_{\text{non-cen}}$ $=\frac{1}{N}\sum_{i=1}^{N}\hat{\gamma}_{\text{non-cen,i}}-\bar{\gamma},$
and the coverages are estimated with
$\widehat{\text{Coverage}}_{\text{cen}}$ $=\frac{1}{N}\sum_{i=1}^{N}I\left\{\hat{\gamma}_{\text{cen,i}}-1.96\hat{se}(\gamma)_{\text{cen,i}}\leq\bar{\gamma}\leq\hat{\gamma}_{\text{cen,i}}+1.96\hat{se}(\gamma)_{\text{cen,i}}\right\},$
$\widehat{\text{Coverage}}_{\text{non-cen}}$ $=\frac{1}{N}\sum_{i=1}^{N}I\left\{\hat{\gamma}_{\text{non-cen,i}}-1.96\hat{se}(\gamma)_{\text{non-cen,i}}\leq\bar{\gamma}\leq\hat{\gamma}_{\text{non-cen,i}}+1.96\hat{se}(\gamma)_{\text{non-cen,i}}\right\}.$
We adopt the bootstrapping method to obtain the $\hat{se}(\gamma)_{\text{cen,i}}$ and $\hat{se}(\gamma)_{\text{non-cen,i}}$ with $B=1000$ sampling for each replicate. We simulate 1000 replicates for each case.

### v2

@@ 3.2 Censoring Weights: Inverse Probability of Censoring Weight
In longitudinal studies, participants are lost to follow-up for various reasons that may be related to the treatment or competing events. This loss of follow-up can result in systematic differences between those who remain in the study and those who do not across treatment groups. While RCTs generally eliminate confounding through randomization, censoring can still introduce bias if the probability of being censored is related to the outcome the participant would have experienced without censoring. To address this issue, censoring weights are applied during the estimation process, aiming to recreate a hypothetical population that would have been observed without censoring. TADA applies inverse probability censoring weights (IPCW) to account for bias resulting from censoring for the time-to-event outcome in the source population.
For each individual, the IPCW is time-varying. Specifically, at each observed event or censoring time point $t$, we estimate:
$w_{i}^{c}(t)\;=\;\frac{1}{\Pr\bigl{(}\Delta_{i}(t)=1\mid\mathbf{X}_{i},A_{i},T_{i}\geq t\bigr{)}},$
where $\Delta_{i}(t)=1$ denotes not censored for individual $i$ at time $t$. These weights are calculated at each distinct event time during follow-up.
We use the Cox PH model 34, 35 to model the hazard of censoring conditional on covariates:
$h_{c}(t\mid A_{i},\textbf{X}_{i})=h_{c0}(t)\cdot\exp\bigl{\{}\eta_{i}(t)\bigr{\}},$
where $h_{c}(t\mid A_{i},\textbf{X}_{i})$ is the hazard function for censoring at time $t$ for individual $i$, $h_{c0}(t)$ is the time-varying baseline hazard function, and $\eta_{i}(t)=\beta_{c}^{\top}(A_{i},\textbf{X}_{i})$ is the linear predictor where $\beta_{c}$ is the vector of regression coefficients and $(A_{i},\textbf{X}_{i})$ is the vector of treatment and covariates.
We estimate the baseline cumulative hazard function $H_{\text{c0}}(t)$ using the Breslow estimator 36:
$H_{\text{c0}}(t)=\sum_{t_{j}\leq t}\frac{d_{j}}{\sum_{l\in R(t_{j})}\exp\bigl{\{}\eta_{i}(t)\bigr{\}}},$
The censoring weight $w_{i}^{c}(t)$ for individual $i$ at time $t$ is defined as the inverse of their survival probability at that time:
$w^{\textbf{c}}_{i}(t)=\frac{1}{S_{c}\bigl{(}t\mid A_{i},\textbf{X}_{i}\bigr{)}},$
where $S_{c}\bigl{(}t\mid A_{i},\textbf{X}_{i}\bigr{)}=\exp\bigl{\{}-H_{c0}(t)\cdot\exp\bigl{(}\eta_{i}(t)\bigr{)}\bigr{\}}$ represents the uncensoring probability at time $t$ given covariates and treatment.
3.3 Final Weights: A Two-staged Scheme
We introduce the final weight for each individual $i$ at time $t$ as the product of the participation weight $w_{i}^{\mathrm{MoM}}(t)$ and the censoring weight $w_{i}^{\mathbf{c}}(t)$, ensuring both population matching and bias correction for censoring as:
$w_{i}^{\textbf{final}}(t)\;=\;w_{i}^{\textbf{MoM}}(t)\times w_{i}^{\textbf{c}}(t).$
When the distribution of the weights is highly skewed, it may be appropriate to consider trimming weights above a certain percentile (e.g., the $95^{\text{th}}$ or $99^{\text{th}}$). Such decisions should be guided by diagnostic assessments, including the inspection of weight distribution and sensitivity analyses across different truncation thresholds, to balance variance reduction against the risk of introducing bias. In the following simulation study introduced in Section 4, we apply an overall $95^{\text{th}}$ percentile trimming threshold after checking the distribution histogram of raw final weights in various settings to achieve the optimal performance for most of scenarios. A systematic sensitivity analysis about various final weights truncation strategies is also conducted and results are presented in Section 6.
This combined weighting approach ensures unbiased treatment effect estimates concerning both censoring and population differences. The weighted source population should closely resemble the target population, demonstrating population balance and supporting the treatment effect estimates. We provide a theoretical justification of unbiased estimation of TADA in section S2 of the Supplementary Materials.
In TADA, we employ a marginal structural model (MSM) to estimate the causal effect of treatment on survival outcomes. Since the weights used in TADA depend solely on baseline covariates and do not involve any time-varying components, we adopt a simplified MSM suitable for this context 37, 38, while TADA remains agnostic to the specific outcome model or assumptions. The MSM is defined as a marginal Cox PH model 34:
$\lambda_{a}(t\mid A_{i})=h_{0}(t)\exp(\beta A_{i}),$
where $\lambda_{a}(t\mid A_{i})$ is the hazard function at time $t$ for individual $i$ with treatment assignment $A_{i}$, $h_{0}(t)$ is the baseline hazard function, and $\beta$ represents the log hazard ratio of the treatment effect.
We estimate $\beta$ using a weighted partial likelihood approach, with the final weights $w_{i}^{\textbf{final}}(t)$ obtained from the two-stage scheme as detailed in Sections 3.1 to 3.3. Applying these final weights creates a pseudo-population in which: (i) the participation weights balance effect modifiers between the source and target populations, ensuring transportability of treatment effects; (ii) the censoring weights reduce bias from incomplete follow-up by assigning greater weights to individuals with higher risk of dropout, approximating the outcomes that would have been observed without censoring under Assumption 3; (iii) treatment is independent of baseline covariates. This framework enables us to estimate the transported treatment effect on the target population using only aggregate-level data.
The weighted partial likelihood function for the Cox model 39 is given by:
$L(\beta)=\prod_{i=1}^{N_{\textit{s}}}\left[\frac{\exp(\beta A_{i})}{\sum_{j\in R(t_{i})}w_{j}^{\textbf{final}}(t_{i})\exp(\beta A_{j})}\right]^{\Delta_{i}w_{i}^{\textbf{final}}(t_{i})},$
where $N_{\textit{s}}$ is the number of individuals in the source sample, $\Delta_{i}$ is the censoring indicator for individual $i$, and $R(t_{i})$ is the risk set of individual $i$ at time $t$. Maximizing this weighted likelihood function provides an estimate of $\beta$, the log hazard ratio, which quantifies the causal effect of the treatment in the target population. We are estimating the variance of $\beta$ using a nonparametric bootstrap method.
4 Simulation Study
The simulation study applies the ADEMP framework proposed by Morris et al. 40 to comprehensively evaluate the effectiveness of TADA in transportability analysis using AgD for a target population, specifically examining its performance in adjusting for censoring. Our primary goal (Aim) is to assess the estimated causal effects of treatment on overall survival across different RCT scenarios, with and without censoring adjustments made via TADA. Detailed descriptions of the Data-generating mechanisms, Estimands, Methods, and Performance measures are provided in the following subsections.
4.1 Data-Generating Mechanism
In each simulation replicate, we generate three datasets named studyIPD, targetAgD, and targetPseudoIPD respectively: studyIPD contains simulated individual-level data from the source population, including treatment status, baseline covariates, event times, censoring times, observed times, and indicators for observed and censored events. targetAgD provides aggregate summary statistics (such as proportions, means, and standard deviations of each baseline covariate) for the target population, matching those in studyData. targetPseudoIPD consists of pseudo individual-level data for the target population, including baseline covariates, treatment status, and event times. We generate both event and censoring times from Weibull distributions, implying that the baseline hazard functions are time-varying 41, 42. Specifically, the Weibull hazard function is defined as:
@@ 4.1 Data-Generating Mechanism
where $\lambda$ is the scale parameter, $\alpha$ is the shape parameter, and $t$ represents time. If $\alpha>1$, the hazard increases over time; if $\alpha<1$, the hazard decreases. Correspondingly, the cumulative hazard and survival functions are:
$H(t)=\lambda t^{\alpha},\quad S(t)=\exp\left\{-H(t)\right\}=\exp\left(-\lambda t^{\alpha}\right).$
@@
To generate studyIPD, we first simulate a source population of 50,000 individuals and then randomly sample 200 individuals for the final dataset. Each individual has three baseline covariates: $X_{1}$ and $X_{2}$ are binary variables generated from binomial distributions with probabilities 0.45 and 0.65, respectively. $X_{3}$ is a continuous variable generated from a standard normal distribution (mean 0, variance 1). Treatment status is assigned randomly to simulate an RCT scenario, with an allocation rate of 1:1.
@@
Event times are generated from a Weibull distribution, incorporating the effects of covariates, treatment and the interaction between treatment and each covariate through the linear predictor $\eta_{\text{event},i}$.
@@
Event times $T_{i}$ are simulated using a Cox PH model with a baseline hazards function following a Weibull distribution as:
@@
In studyIPD, we generate censoring times for each individual similarly to event times. A linear predictor $\eta_{\text{censor},i}$ is constructed for each individual to incorporate the effects of covariates on the hazard function.
@@
where $\beta_{X_{1}}^{c}=-0.2$, $\beta_{X_{2}}^{c}=0.4$ and $\beta_{X_{3}}^{c}=-0.1$ as the main effect of baseline covariates, $\beta_{\text{trt}}^{c}=0.25$ as the treatment main effect and $\beta_{0}$ as the intercept. We choose various candidates of $\beta_{0}$ to directly adjust the overall censoring proportion among studyIPD. We set $\beta_{0}$ as 2.5, 3.3, 3.7 and 4.3 respectively to obtain studyIPD with around 20%, 30%, 40% and 50% overall censoring proportion while keeping other coefficients being consistent.
@@
For each individual, the observed time $U_{i}$ is determined by:
@@
The censoring indicator $\Delta_{i}$ is defined as:
@@
Similarly, we generate targetAgD by simulating a larger target population of 200,000 individuals and then randomly sampling 1,000 individuals for the final dataset. Each individual has the same three baseline covariates as in the source population ($X_{1}$, $X_{2}$, and $X_{3}$) with different distribution characteristics: $X_{1}$ is binary and generated from a binomial distribution with probability 0.35. $X_{2}$ is binary and generated from a binomial distribution with probability 0.55. $X_{3}$ is continuous and generated from a normal distribution with a mean of 0.21 and a variance of 1.5. Since targetAgD is an aggregate-level dataset, we summarize the proportions of $X_{1}$ and $X_{2}$, and the mean and standard deviation of $X_{3}$ as the final data.
The individual-level covariate data for targetPseudoIPD are generated concurrently before aggregation. To obtain the “true” treatment effect in the target population for benchmarking, we assign treatment status and generate event times for each individual in targetPseudoIPD. Treatment status is randomly assigned with a 50% probability to simulate an RCT scenario. We use the same Weibull distribution parameters ($\alpha_{\text{event}}$, $\lambda_{\text{event}}$), the same linear predictor structure $\eta_{\text{event},i}$ and coefficients as in studyIPD, applying them to targetPseudoIPD to ensure consistency.
4.2 Estimands, Targets, and Methods
@@
  • Adjusting censoring: In this case, TADA balances the source and target population with the participation weights introduced in Section 3.1, as well as adjusting the bias resulting from censoring via IPCW introduced in Section 3.2. The final weights are the product of the two as Section 3.3.
  • Ignoring censoring: In this case, TADA ignores the impact of censoring and only balances the source and target population with the participation weights introduced in Section 3.1. The final weights are equivalent to the participation weights estimated by the method of moments.
For both strategies, we estimate the average marginal hazard ratios using a Cox PH model based on multiple replicates. We apply the Cox PH model on targetPseudoIPD to estimate average marginal hazard ratios as the pseudo “true HR” of the target population as the comparison benchmark.
@@ 4.3 Performance Measures
The performance measures of interest are the bias of the estimators calculated using the above methods and their coverage of 95% confidence intervals compared to the analysis in the IPD target data. Let $\hat{\gamma}_{\text{cen,r}}$ be the estimated marginal hazard ratios for the $r=1,2,\cdots,R$-th simulated replicate when considering the adjustment of censoring. Let $\hat{\gamma}_{\text{non-cen,r}}$ be the estimated marginal hazard ratios for the $r=1,2,\cdots,R$-th simulated replicate when ignoring the adjustment of censoring. With the pseudo “true HR” of target population being $\overline{\gamma}$, the bias of two cases aforementioned are estimated with
$\widehat{\text{Bias}}_{\text{cen}}=\frac{1}{R}\sum_{r=1}^{R}\left(\hat{\gamma}_{\text{cen,r}}-\overline{\gamma}\right),\quad\widehat{\text{Bias}}_{\text{non-cen}}=\frac{1}{R}\sum_{r=1}^{R}\left(\hat{\gamma}_{\text{non-cen,r}}-\overline{\gamma}\right).$
For each replicate $r=1,\ldots,R$, we construct a 95% percentile-based bootstrap confidence interval $\bigl{[}\widehat{\mathrm{CI}}_{\mathrm{low},r},\;\widehat{\mathrm{CI}}_{\mathrm{high},r}\bigr{]}$ for the hazard ratio, where
$\widehat{\mathrm{CI}}_{\mathrm{low},r}=\text{Quantile}_{0.025}\bigl{(}\hat{\gamma}_{r}^{*(1)},\dots,\hat{\gamma}_{r}^{*(B)}\bigr{)},\quad\widehat{\mathrm{CI}}_{\mathrm{high},r}=\text{Quantile}_{0.975}\bigl{(}\hat{\gamma}_{r}^{*(1)},\dots,\hat{\gamma}_{r}^{*(B)}\bigr{)}.$
We then define the coverage as the proportion of replicates for which $\overline{\gamma}$ lies within this interval:
$\widehat{\mathrm{Coverage}}_{\mathrm{cen}}=\frac{1}{R}\sum_{r=1}^{R}I\Bigl{\{}\overline{\gamma}\in\bigl{[}\widehat{\mathrm{CI}}_{\mathrm{low},r}^{(\mathrm{cen})},\widehat{\mathrm{CI}}_{\mathrm{high},r}^{(\mathrm{cen})}\bigr{]}\Bigr{\}},\quad\widehat{\mathrm{Coverage}}_{\mathrm{non\text{-}cen}}=\frac{1}{R}\sum_{r=1}^{R}I\Bigl{\{}\overline{\gamma}\in\bigl{[}\widehat{\mathrm{CI}}_{\mathrm{low},r}^{(\mathrm{non\text{-}cen})},\widehat{\mathrm{CI}}_{\mathrm{high},r}^{(\mathrm{non\text{-}cen})}\bigr{]}\Bigr{\}}.$
Considering the trade-off between accuracy and computational cost in the presence of time-varying weight estimation in our simulation, we simulate $R$ = 500 replicates for each scenario and within each replicate generate $B=200$ bootstrap resamples.